Calculate order and frequently

You must show all work (i.e., step-by-step computation) and label the final answer.

  1. An air-conditioning manufacturing plant has a vendor next door supplying fan blades almost instantaneously on order. The demand is 1,000 blades/week. The ordering cost for blades is $7/order and the purchase price for each blade is $12. The nominal (annual) interest rate assessed to the investment for this purpose is 10%. The plant does not want to lose any customer. How many blades should the manager of the plant order at minimal yearly cost and how frequently should it be ordered?
Answer #1

Weekly demand = 1000

Annual demand (D) = Weekly Demand * Months per year * Weeks per month = 48,000 (1000*12*4)

Ordering Cost(O) = $7

Holding Cost per year (H) = $1.2 ($12*10%)

Optimal order size = sqrt((2*D*O)/H

= sqrt(2*48000*7)/(1.2)

Optimal order size = 748.331477or 748 (approx)

Order frequency = 365/number of orders

= 365/(D/optimal order size)

= 365/(48,000/748.331477)

= 365/64.1426981

= 5.69043727 or 6 days


answered by: anonymous